Given that IS is Y = 2500-50i, and the interest rate reaction function is ip = 2 + 0.5(P-P*), where ip is the central bank policy rate. Assume that i (interest rate) is 0.5 + ip (interest rate is a wedge over the policy rate). Determine the AD function. If AS function was Y = 0.25(P-P*) + 2000 determine Y* and P* and inflation rate. Assume P* = 105.
1
Expert's answer
2020-05-18T11:51:04-0400

"i=0.5+2+0.5(P-P^*)=2.5+0.5(P-P^*)"

"Y=2500-50(2.5+0.5(P\u2212P ^\u2217\n ))=2500-125-25(P\u2212P ^\u2217\n )=2375-25(P\u2212P^\u2217\n )"

"AD = Y =2375-25(P\u2212P^\u2217)"

Y*-equlibrium output

If P*=105

"AD=AS=Y^*"

"2375-25(P-105)=0.25(P-105)+2000"

"25.25(P-105)=375"

"P-105=14.85"

"P=119.85"

"Y^*=2375-25*14.85=2003.75"

ir-inflation rate. 

"ir=\frac{P-P^*}{P^*}*100=\frac{14.85}{105}*100=14.14"

ir=14.14%